# Mechanistic Modeling and Analysis of Transport Phenomena for Continuous Lyophilization

### v2

@@ Conclusion
Data Availability
All software and data will be made available upon publication of the manuscript. Both MATLAB and Julia implementations will be provided.
@@ Appendix A Analysis of transport phenomena
Biot number and spatial dimensions
As discussed in the main text, the freezing step is modeled using the lumped capacity method, whereas the drying steps are analyzed in one spatial dimension, a vertical direction. Various studies have shown via simulation that variation in the radial direction is negligible [17, 13]. This section provides a detailed analysis from the theoretical perspective to justify this assumption and briefly discusses strategies to modify the model for situations where this assumption is not true.
@@
where $L$ is the characteristic length or length scale, $h$ is the heat transfer coefficient, and $k$ is the thermal conductivity. In general, the length scale is defined as the ratio of the volume to the surface area, which results in $r_{\textrm{o}}/2$ for a long cylinder. However, for conservative analysis, it is recommended to use the actual radius $r_{\textrm{o}}$ instead [45]. The Biot number is defined as the ratio of the internal conduction resistance to the external convection resistance, and so a small Bi value implies that the resulting temperature gradient should also be small [31]. In various heat transfer textbooks, the value of 0.1 is recommended as a guideline; if Bi $<0.1$, a lumped capacity model can be used. However, in practice, a lumped capacity model could still be valid even for Bi $>0.1$, depending on the system, which is analyzed here.
@@
Finally, the errors (differences) between both models for different values of Bi are shown in Fig. A.3. Of course, if the Biot is too high, the lumped capacity model will not be accurate. That situation could happen, e.g., when a very large vial is used or the heat transfer coefficient is much higher than usual. In such cases, a high-fidelity model might be needed [13]. However, the analysis in this section suggests that a lumped capacity model is highly accurate for typical lyophilization applications, in particular for unit doses.
@@
Mass transport in the dried product
In secondary drying, the two important mechanisms for mass transport are (1) vapor phase or water vapor transport through the porous dried product and (2) desorption of bound water from the solid surface of the dried product. Previous simulation and experimental data suggest that desorption is the rate-limiting step [37, 23, 4, 24, 25, 36], and so the equations/terms related to vapor phase transport are typically omitted from recently published models, including our model. This section provides an analysis from the perspective of transport phenomena to justify this assumption.
First, consider the vapor phase transport through the porous dried product. Modeling the transport of liquids and gases in porous media is challenging as it could entail a variety of mechanisms besides ordinary diffusion (e.g., Knudsen diffusion), and thus no fully satisfactory theory currently exists [46]. Nevertheless, relative success has been achieved by using the concept of an effective diffusivity; in this analysis, we follow the procedure described in Ref. [31]. Define the effective diffusivity $D_{\textrm{e}}$ by
@@
With the effective diffusivity calculated by (A.2)–(A.5), the total mass flux can be calculated using Fick’s law of diffusion, with the binary diffusion coefficient replaced by the effective diffusivity. Then, a continuity (mass balance) equation for the water vapor can be written and coupled to the mechanistic model proposed in the main paper. The key benefit of this technique is that the implementation closely resembles the approach for solving problems entailing ordinary diffusion of a binary mixture, which is relatively straightforward. Nevertheless, additional information is needed, including the porosity, tortuosity, and average pore radius. A more complicated strategy that is used in the literature is to apply the dusty gas model, which is derived from the Maxwell-Stefan Equation (see examples in [16, 19]). This approach, however, requires many more parameters and equations.
Next, we compare the significance of vapor phase transport and desorption. Scale analysis can be used to analyze the contribution of each transport process (see various examples in [31]). Properly estimating the time scale of each phenomenon can help identify the key process that should be considered in the model and omit unimportant phenomena. The time scale (aka time constant) for mass diffusion can be approximated by [47, 31]
@@
where $t_{\text{des}}$ is the time scale for desorption and $k_{\textrm{d}}$ is the desorption constant defined in (43). By comparing $t_{\text{dff}}$ and $t_{\text{des}}$, the limiting step for mass transfer in the secondary drying step can be identified, without any simulation or experimental data.
For numerical comparison, a set of parameter values are obtained from Refs. [15, 17], namely $\bar{r}=5$$\times$$10^{-6}$ m, $\epsilon=0.815$, $\tau=1.2$, and $k_{\textrm{d}}=7.8$$\times$$10^{-5}$ s-1. The binary coefficient of water in air is obtained from Ref. [31], which is $D_{\textrm{g}}=1.97$$\times$$10^{-5}$ m2/s at 256 K and 1 atm. In secondary drying, the operating pressure is much lower than 1 atm, and so the actual $D_{\textrm{g}}$ is lower as the diffusion coefficient of gas mixtures at low pressure is inversely proportional to the pressure [31, 46], which makes our analysis more conservative. The length scale for this analysis is set to 0.01 m, a typical vial/sample size (thickness/diameter). Substituting these values into (A.6) and (A.7) gives $t_{\text{dff}}=7.5$ s and $t_{\text{des}}=3.6$ h. Obviously, the time scale for diffusion is much smaller than for desorption by several orders of magnitude, indicating that desorption is the rate-limiting step. A typical drying time of secondary drying is on the order of several hours, agreeing with the calculated $t_{\text{des}}$ of about 3.6 h. On the other hand, the time scale for diffusion is only 7.5 s, meaning that diffusion is much faster and so can be safely omitted from the model. Even some of the parameters are changed by an order of magnitude, this conclusion does not change.***This analysis holds for lyophilization of unit doses only; for lyophilization of bulk material (e.g., in food applications), the length scale of the product could be much larger, and so the conclusion might be different.
Finally, there is another mass transfer mechanism that is not commonly discussed or modeled, namely diffusion of water in the solid matrix (not in the pores), i.e., solid-phase diffusion [37]. Diffusion in the solid matrix is very slow, and so it could limit the mass transport in secondary drying. Modeling this phenomenon is challenging as it requires detailed knowledge of the solid matrix, particle size, and locations of bound water in the matrix, which is generally not available. A similar approximation to (A.6) can be carried out. By using the reported length scale of about $5$$\times$$10^{-7}$ m and diffusion coefficient of water in the solid phase of about $7$$\times$$10^{-16}$ m2/s [37], the time scale for solid-phase diffusion is about 10 min, which is still much lower than for desorption.
To summarize, this section details the scale analysis that can be efficiently used to evaluate the effects of various mass transfer mechanisms. For lyophilization of pharmaceuticals in unit doses, our analysis shows that desorption of bound water is the rate-limiting step for mass transfer, and so other mechanisms can be omitted from the model, agreeing with both simulation and experimental data in the literature.
@@ Appendix B Supplemental equations
In Section 3.2.1, if the total liquid volume $V_{\textrm{l}}$ and mass fraction of a solute $x_{\textrm{s}}$ are reported, the density of the liquid phase $\rho_{\textrm{l}}$ is
@@
Then (1) can be used as usual.
@@
For the primary drying model (Section 3.3), the volume of the frozen region is
@@
For the secondary drying model (Section 3.4), the volume of the dried region is
@@
Appendix C Numerical methods
This appendix comprehensively describes numerical techniques required for solving the model equations efficiently. With the given procedure, our model can be simulated accurately within 1 s on a normal laptop.
For the freezing model, since the lumped capacity approximation is used, the resulting equations are ordinary differential equations (ODEs), and so the equations can be integrated directly with a proper ODE solver. This work uses MATLAB’s ode15s.
For PDEs in the primary drying and secondary drying models, the method of lines [48] is recommended. The method of lines consists of two steps. First, the PDEs are discretized spatially to produce a system of ODEs. Second, the resulting ODEs are integrated with a proper ODE solver (MATLAB’s ode15s in this work).
For the primary drying model, first consider the PDE (32). As the equation involves a moving interface $S(t)$, define a new variable
$\xi=\frac{z-S}{H-S},$ (C.1)
where $\xi$ is the dimensionless position with respect to the moving interface (aka sublimation front). This dimensionless position varies from 0 to 1, in which $\xi=0$ at $z=S$ and $\xi=1$ at $z=H$. Consequently, the PDE (32) and boundary conditions (38) and (39) become
$\rho_{\textrm{f}}\,C_{p,\textrm{f}}\,\dfrac{\partial T}{\partial t}=\frac{k_{\textrm{f}}}{(H-S)^{2}}\dfrac{\partial^{2}T}{\partial\xi^{2}}-\frac{\rho_{\textrm{f}}\,C_{p,\textrm{f}}\,(\xi-1)}{H-S}\frac{dS}{dt}\frac{\partial T}{\partial\xi}+\frac{Q_{\textrm{rad}}}{V_{\textrm{f}}},\qquad 0<\xi<1,$ (C.2)
$-\frac{k_{\textrm{f}}}{H-S}\dfrac{\partial T}{\partial\xi}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad\xi=1,$ (C.3)
$N_{\textrm{w}}\Delta H_{\textrm{sub}}=\frac{k_{\textrm{f}}}{H-S}\dfrac{\partial T}{\partial\xi}+\sigma\mathcal{F}_{\textrm{s1}}(T_{\textrm{u}}^{4}-T^{4}),\qquad\xi=0.$ (C.4)
To spatially discretize the PDE, define
$\Delta\xi=\frac{1}{n_{z}-1},$ (C.5)
$\xi=(j-1)\Delta\xi,$ (C.6)
where $n_{z}$ is the number of grid points, $\Delta\xi$ is the distance between each grid point, and $j$ is the integer index for discretization, in which $j=1$ at $z=S$ and $j=n_{z}$ at $z=H$. Discretizing (C.2)–(C.4) using the second-order finite difference scheme results in
$\begin{split}\dfrac{dT_{j}}{dt}=\frac{k_{\textrm{f}}}{\rho_{\textrm{f}}C_{p,\textrm{f}}(H-S)^{2}}\!\left(\dfrac{T_{j+1}-2T_{j}+T_{j-1}}{\Delta\xi^{2}}\right)\!-\frac{j\Delta\xi-1}{H-S}\frac{dS}{dt}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!+\frac{Q_{\textrm{rad}}}{\rho_{\textrm{f}}C_{p,\textrm{f}}V_{\textrm{f}}},\\ \qquad j=1,2,\dots,n_{z},\end{split}$ (C.7)
$-\frac{k_{\textrm{f}}}{H-S}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!=h_{\textrm{b}}(T_{j}-T_{\textrm{b}}),\qquad j=n_{z},$ (C.8)
$N_{\textrm{w}}\Delta H_{\textrm{sub}}=\frac{k_{\textrm{f}}}{H-S}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!+\sigma\mathcal{F}_{\textrm{s1}}\!\left(T_{\textrm{u}}^{4}-T_{j}^{4}\right),\qquad j=1,$ (C.9)
where $T_{j}$ is the product temperature at position $j$. The ghost point at $j=n_{z}+1$ in (C.8) can be eliminated by substituting $T_{j=n_{z}+1}$ into the governing equation (C.7) for $j=n_{z}$. Similarly, $T_{j=0}$ in (C.9) can be eliminated by substituting $T_{j=0}$ into the governing equation (C.7) for $j=1$. The above nondimensionalization and discretization techniques result in a system of nonlinear ODEs defined on a moving-grid system.
$\Delta z=\frac{H}{n_{z}-1},$ (C.10)
$z=(j-1)\Delta z,$ (C.11)
where $\Delta z$ is the distance between each grid point, $j=1$ at $z=0$, and $j=n_{z}$ at $z=H$. Discretizing (42), (43), (45), and (46) using the second-order finite difference scheme results in
$\frac{dT_{j}}{dt}=\frac{k_{\textrm{e}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}}\!\left(\frac{T_{j+1}-2T_{j}+T_{j-1}}{\Delta z^{2}}\right)\!+\frac{\rho_{d}\Delta H_{\textrm{des}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}}\!\left(\dfrac{dc_{\textrm{w},j}}{dt}\right)\!+\frac{Q_{\textrm{rad}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}V_{\textrm{e}}},\qquad j=1,2,\dots,n_{z},$ (C.12)
$\dfrac{dc_{\textrm{w},j}}{dt}=-f_{\textrm{a}}e^{-E_{\textrm{a}}/RT_{j}}c_{\textrm{w},j},\qquad j=1,2,\dots,n_{z},$ (C.13)
$-k_{\textrm{e}}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta z}\right)\!=h_{\textrm{b}}(T_{j}-T_{\textrm{b}}),\qquad j=n_{z},$ (C.14)
$-k_{\textrm{e}}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta z}\right)\!=\sigma\mathcal{F}_{\textrm{s1}}\!\left(T_{\textrm{u}}^{4}-T_{j}^{4}\right),\qquad j=1,$ (C.15)

### v3

@@ 6 Conclusion
Data and Code Availability
Software and data used in this work are available at https://github.com/PrakitrSrisuma/ContLyo-modeling. All simulations, calculations, and results presented in this article were performed and generated using MATLAB, but an equivalent Julia implementation of the code is also provided.
@@ Appendix A Analysis of Transport Phenomena
A.1 Biot Number and Spatial Dimensions
As discussed in the main text, the freezing step is modeled using the lumped capacity method, whereas the drying steps are analyzed in one spatial dimension, a vertical direction. Various studies have shown via simulation that variation in the radial direction is negligible [Sheehan1998Modeling, Deck2024Freezing2D]. This section provides a detailed analysis from the theoretical perspective to justify this assumption and briefly discusses strategies to modify the model for situations where this assumption is not true.
@@ A.1 Biot Number and Spatial Dimensions
where $L$ is the characteristic length or length scale, $h$ is the heat transfer coefficient, and $k$ is the thermal conductivity. In general, the length scale is defined as the ratio of the volume to the surface area, which results in $r_{\textrm{o}}/2$ for a long cylinder. However, for conservative analysis, it is recommended to use the actual radius $r_{\textrm{o}}$ instead [Incropera2007HeatTransfer]. The Biot number is defined as the ratio of the internal conduction resistance to the external convection resistance, and so a small Bi value implies that the resulting temperature gradient should also be small [Mills1995HeatTransfer]. In various heat transfer textbooks, the value of 0.1 is recommended as a guideline; if Bi <0.1<0.1, a lumped capacity model can be used. However, in practice, a lumped capacity model could still be valid even for Bi >0.1>0.1, depending on the system, which is analyzed here.
@@
Finally, the errors (differences) between both models for different values of Bi are shown in Figure A.3. Of course, if the Biot is too high, the lumped capacity model will not be accurate. That situation could happen, e.g., when a very large vial is used or the heat transfer coefficient is much higher than usual. In such cases, a high-fidelity model might be needed [Deck2024Freezing2D]. However, the analysis in this section suggests that a lumped capacity model is highly accurate for typical lyophilization applications, in particular for unit doses.
@@
A.2 Mass Transport in the Dried Product
In secondary drying, the two important mechanisms for mass transport are (1) vapor phase or water vapor transport through the porous dried product and (2) desorption of bound water from the solid surface of the dried product. Previous simulation and experimental data suggest that desorption is the rate-limiting step [Pikal1990SecDrying_Kinetics, Fissore2011SecDryingMonitor, Fissore2015Review, Sahni2017Simplified, Yoon2021Sec0D1D3D, Srisuma2024Obs], and so the equations/terms related to vapor phase transport are typically omitted from recently published models, including our model. This section provides an analysis from the perspective of transport phenomena to justify this assumption.
First, consider the vapor phase transport through the porous dried product. Modeling the transport of liquids and gases in porous media is challenging as it could entail a variety of mechanisms besides ordinary diffusion (e.g., Knudsen diffusion), and thus no fully satisfactory theory currently exists [Bird2002Transport]. Nevertheless, relative success has been achieved by using the concept of an effective diffusivity; in this analysis, we follow the procedure described in [Mills1995HeatTransfer]. Define the effective diffusivity $D_{\textrm{e}}$ by
@@ A.2 Mass Transport in the Dried Product
With the effective diffusivity calculated by Equations (A.2)–(A.5), the total mass flux can be calculated using Fick’s law of diffusion, with the binary diffusion coefficient replaced by the effective diffusivity. Then, a continuity (mass balance) equation for the water vapor can be written and coupled to the mechanistic model proposed in the main paper. The key benefit of this technique is that the implementation closely resembles the approach for solving problems entailing ordinary diffusion of a binary mixture, which is relatively straightforward. Nevertheless, additional information is needed, including the porosity, tortuosity, and average pore radius. A more complicated strategy that is used in the literature is to apply the dusty gas model, which is derived from the Maxwell-Stefan Equation (see examples in [Sadikoglu1997Modeling, Veraldi2008SimBatchModels]). This approach, however, requires many more parameters and equations.
Next, we compare the significance of vapor phase transport and desorption. Scale analysis can be used to analyze the contribution of each transport process (see various examples in [Mills1995HeatTransfer]). Properly estimating the time scale of each phenomenon can help identify the key process that should be considered in the model and omit unimportant phenomena. The time scale (aka time constant) for mass diffusion can be approximated by [Crank1979Diffusion, Mills1995HeatTransfer]
@@
where $t_{\text{des}}$ is the time scale for desorption and $k_{\textrm{d}}$ is the desorption constant defined in Equation (43). By comparing $t_{\text{dff}}$ and $t_{\text{des}}$, the limiting step for mass transfer in the secondary drying step can be identified, without any simulation or experimental data.
For numerical comparison, a set of parameter values are obtained from [Liapis1994Original, Sheehan1998Modeling], namely $\bar{r}=5$$\times$$10^{-6}$ m, $\epsilon=0.815$, $\tau=1.2$, and $k_{\textrm{d}}=7.8$$\times$$10^{-5}$ s-1. The binary coefficient of water in air is obtained from [Mills1995HeatTransfer], which is $D_{\textrm{g}}=1.97$$\times$$10^{-5}$ m2/s at 256 K and 1 atm. In secondary drying, the operating pressure is much lower than 1 atm, and so the actual $D_{\textrm{g}}$ is lower as the diffusion coefficient of gas mixtures at low pressure is inversely proportional to the pressure [Mills1995HeatTransfer, Bird2002Transport], which makes our analysis more conservative. The length scale for this analysis is set to 0.01 m, a typical vial/sample size (thickness/diameter). Substituting these values into Equations (A.6) and (A.7) gives $t_{\text{dff}}=7.5$ s and $t_{\text{des}}=3.6$ h. Obviously, the time scale for diffusion is much smaller than for desorption by several orders of magnitude, indicating that desorption is the rate-limiting step. A typical drying time of secondary drying is on the order of several hours, agreeing with the calculated $t_{\text{des}}$ of about 3.6 h. On the other hand, the time scale for diffusion is only 7.5 s, meaning that diffusion is much faster and so can be safely omitted from the model. Even some of the parameters are changed by an order of magnitude, this conclusion does not change. [^2] This analysis holds for freeze drying of unit doses only; for freeze drying of bulk material (e.g., in food applications), the length scale of the product could be much larger, and so the conclusion might be different.
[…]
To summarize, this section details the scale analysis that can be efficiently used to evaluate the effects of various mass transfer mechanisms. For freeze drying of (bio)pharmaceuticals in unit doses, our analysis shows that desorption of bound water is the rate-limiting step for mass transfer, and so other mechanisms can be omitted from the model, agreeing with both simulation and experimental data in the literature.
@@ Appendix B Supplemental Equations
For the freezing model, if the total liquid volume $V_{\textrm{l}}$ and mass fraction of a solute $x_{\textrm{s}}$ are reported, the density of the liquid phase $\rho_{\textrm{l}}$ is
@@
Then, Equation (1) can be used as usual.
@@
For the primary drying model, the volume of the frozen region is
@@
For the secondary drying model, the volume of the dried region is
@@
This article presents the first mechanistic model for continuous lyophilization (aka freeze drying). The model accurately describes the key phenomena and critical process parameters in continuous lyophilization, which can be used for process design, optimization, and control. Results from this work could drive the transition of (bio)pharmaceutical freeze drying from batch processes to continuous manufacturing.